Wtite in standard form the equation of a line containing the points (0,-2) and (4, 5)
And in what grade are they teaching these

Answer by Alan3354(69443)

Wtite in standard form the equation of a line containing the points
(0,-2) and (4,5)
--------------
Find the slope, m
m = diffy/diffx = 7/4
y = mx + b
(0,-2) is the y-intercept
----
y = (7/4)x - 2
7x/4 = y+2
7x = 4y + 8
7x - 4y = 8
============
And in what grade are they teaching these
8th grade, sometimes earlier